OpenStudy (anonymous):

\[\LARGE \lim_{x\to0}\frac{1-\cos 2x}{x\cdot \sin x }\] A) 1 B) 2 C) 3 D) 4 I need a hint please ...

OpenStudy (turingtest):

\[\cos(2x)=1-2\sin^2x\]should help I think

OpenStudy (anonymous):

Thank you... I'll see what I can do :D

OpenStudy (anonymous):

I think I made a mistake somewhere... attempted: \[ \lim_{x\to0}\frac{1-(1-2\sin x)}{x\cdot \sin x}=\lim_{x\to0}2\cdot \frac{\sin x}{x}\cdot \sin x=\] \[=2\cdot \sin 0= 2\cdot 0=0 \] What's wrong ? :(

OpenStudy (turingtest):

you forgot the squared

OpenStudy (anonymous):

I just made a mistake... it is \[\lim_{x\to0}\frac{1-(1-2 \sin^2x)}{x\cdot \sin x}\]

OpenStudy (turingtest):

right, follow that through

OpenStudy (anonymous):

but still the answer holds , what's wrong?

OpenStudy (turingtest):

no you made a mistake how did you wind up with sin^2 in the numerator and no sin in the denom? you made an algebra mistake

OpenStudy (turingtest):

where did sin in the denominator go?

OpenStudy (turingtest):

...also you can't plug in zero where you did, because you still have x in the denominator.

OpenStudy (anonymous):

so... \[\lim_{x\to0}\frac{1-(1-2 \sin^2x)}{x\cdot \sin x}= \lim_{x\to0}\frac{2\sin x}{x}=2 !!\] Dumb answer !! how it is ??

OpenStudy (anonymous):

I didn't plug nothing .. that's why should be wrong, but I'm confused ^o)

OpenStudy (turingtest):

you are supposed to recognize that\[\lim_{x\to0}\frac{1-(1-2 \sin^2x)}{x\cdot \sin x}= \lim_{x\to0}\frac{2\sin x}{x}=2 \lim_{x\to0}\frac{\sin x}{x}=2(1)=2\]did you not know that\[ \lim_{x\to0}\frac{\sin x}{x}=1\]?

OpenStudy (anonymous):

Absolutely .... I don't need to lie check out my work that I did a few minutes ago, I USED that formula... but here's a question : We didn't plug nothing is it supposed x to attend to zero ?? x->0 we didn't plug anything in

OpenStudy (turingtest):

yes, x tends to 0 it doesn't have to "become" zero in the limit it is a removable discontinuity at x=0 because\[ \lim_{x\to0^-}\frac{\sin x}{x}=\lim_{x\to0^+}\frac{\sin x}{x}=1\]even though the function is not defined at x=0 that is the definition of a removable discontinuity: no plug-ins allowed

OpenStudy (anonymous):

ahhh... now I get it, anyway , without your first hint I would not be able even to start it... Thanks in advance ;)

OpenStudy (turingtest):

you're welcome :)